### v1

@@ 3.4 Significant treatment success: emergence of oscillations
As we anticipated, it is well known that solutions of Eq. (U3) may exhibit persistent oscillations for some parameter values; this is in line with the fact that the non-spatial model of Eq. (2.6) presents a Hopf bifurcation. Fig. 7a shows the bifurcation in the parameters $\alpha$, obtained using the software xppaut, which allows to study the stability of equilibria and limit cycles through numerical continuation. Fig. 7b-c shows the time series for three different values of $\alpha$. We remark that a similar behaviour is also observed, for example, as $\beta$} increases or as $q_{v}$ decreases. The one-dimensional setting is not ideal for studying oscillations, as local tumour extinctions due to stochasticity are very common: indeed, the typical behaviour observed in such parameter ranges is analogous to the one described below in relation to Fig. 11. We therefore restrict our attention to two spatial dimensions: the higher number of cells involved allow us to describe oscillations, as well as stochastic-driven partial and complete eradication.
Figure 7: (a) One parameter bifurcation in $\alpha$ of Eq. (2.2), with other parameters as in Table 1 and $q_{v}=0.167\,$h-1. The green dots show the maximum and minimum values of $U/K$ for the stable limit cycle oscillations. The solid lines show the equilibrium value of $U$ divided by $K$; the line is red if the equilibrium is stable and black if it is unstable. Hopf bifurcations are denoted by HB. Observe that for low values of $\alpha$, the infection-free equilibrium close to carrying capacity is stable. (b-c) Numerical simulation of Eq. (2.5) with the parameters as in Table 1 and different values of the viral burst size $\alpha$. The solid blue lines represent uninfected cells, the solid red lines infected cells and the dotted purple lines represent viral density multiplied by $q_{v}/(\alpha q)$; the two latter quantities superimpose almost exactly. The oscillations become wider as $\alpha$ increases, in accordance with the bifurcation diagram of panel (a).
@@
Fig. 8, along with the video accompanying it (see electronic supplementary material S7), shows an overall excellent quantitative agreement between numerical solutions of the system of PDEs (U3) and the average over five numerical simulations of the corresponding agent-based model in two spatial dimensions. After the initial infection reaches the tumour boundary, the number of cancer cells decreases significantly; as a consequence, stochasticity becomes relevant and, in some regions, cells go extinct. Fig. 8a shows the inhomogeneous pattern of uninfected cells caused by this. As time passes, the spatial inhomogeneities quickly disappear, and both models show similar persistent oscillations in the tumour centre around the equilibrium values (coherently with what the bifurcation diagram shows). A slight delay of the oscillations is the main difference between the two modelling approaches: this is easy to see from the total number of cells, depicted in Fig. 10a. The invasion speed of the tumour in the surrounding area is still $2\sqrt{Dp}$ and the uninfected cell density at the front is approximately given by Eq. (3.2), despite some fluctuations. It is interesting to remark that highly dynamic patterns may emerge when viral injections at multiple sites are considered [4].
Fig. 10a shows the comparison with the solution of Eq. (U2) to remark once again that oscillations cannot be obtained in the two-equations model even for high infection rates. Furthermore, Fig. 10b shows that the oscillations are almost non existent for lower viral diffusion, resembling the behaviour of Eq. (U2): this confirms once again the excellent quality of the quasi-steady assumption as spatial viral dynamics become negligible and the reactions involving the virus increase their rate.
@@
Figure 10: Comparison between the sum of tumour cells in different scenarios. The parameters employed are the ones given in Table 1, $\alpha=3500\;$viruses/cells and $q_{v}=1.67\times 10^{-1}\;$h-1; $D_{v}$ assumes different values in panels (b) and (c). Solid lines represent the average of five agent-based simulations in the case of undirected cell movement (yellow and red) and pressure-driven cell movement (purple and green). All the other lines represent the sum of cells obtained from continuous models. The dashed blue lines refer to the solution of Eq. (U2); they are in a different colour than the other PDE solutions, to highlight the fact that it is related to a two-equations model, which does not exhibit persistent oscillations. The black lines refer to the solutions of either Eq. (U3) or Eq. (P3): in panels (b) and (c) we show for simplicity only the case of lower viral diffusion, as the reference cases are already shown in panel (a) and their qualitative behaviour is similar to the discrete models.
@@
Fig. 9a, along with the video accompanying it (see electronic supplementary material S8), shows that, as in the previous case, the number of cancer cells decreases significantly after the initial infection reaches the tumour boundary in both models and, in some regions, cells go extinct, originating inhomogeneous patterns. However, the subsequent regrowth does not allow a fast spatial spread due to the slow movement that cells experience when the pressure is low. Fig. 9b indeed shows that uninfected cells take a long time in the agent-based model to move from the hollow ring into the central region; on the other hand, in the case of the PDE, the repopulation is much faster: there is never a complete extinction in the centre and the increase of cells’ number in some areas is caused by regrowth rather than movement. This difference is particularly evident from the second peak in Fig. 10b. The difference between the two models tends to disappear as time passes, but Fig. 9c still shows some dissimilarities.
Both models have an invasion speed much lower than $\sqrt{Dp/2}$: indeed, the nonlinear diffusion is affected by the scale of the cell density, as explained in Appendix A. In our case, the oscillatory behaviour of the system does not allow to easily identify a unique, invariant speed. Fig. 9 shows that we can obtain a good approximation by setting $\bar{u}=1800\;$cells/mm2, which corresponds to an average over time of the peak. The change in the speed in presence of effective infections is a significant difference from the case of undirected cell movement and, as a result, the overall cell numbers involved are lower in the second case (see Fig. 10a).
@@
The oscillations analysed so far are able to drive the tumour at low densities, in which stochastic fluctuations may cause the death of all the agents in some regions; nevertheless, oscillations are not wide enough to lead the whole tumour close to extinction. Our goal is now to increase the effectiveness of the therapy and discuss the possibility of a full tumour remission. We therefore decrease the value of $q_{v}$ to half of the reference value, so that the oscillations can reach lower cell densities. We also consider an initial wide viral infection (i.e., $R_{v}=R_{u}$) to study the effect of a higher initial viral load. Fig. 11 summarises the cell number of one hundred simulations for both cell movement at time t=1500t=1500\;h. In both cases, the wide majority of the simulations shows the extinction of the tumour; the case of undirected cell movement appears more successful, with only eight simulations with a surviving tumour (in contrast to nineteen in the other case) and an overall lower cell density at the final time. Only uninfected cells survive until the end of the simulation: indeed, small uninfected population tends to regrow, while the infection requires both viral particle and uninfected cells in the same locations to survive.
@@
Figure 11: (a)-(b) Violin plots of the final cell number at time t=1500t=1500\;h obtained from one hundred simulations of the agent-based models with the parameters given in Table 1, $\alpha=3500\;$viruses/cells and $q_{v}=8.33\times 10^{-2}\;$h-1; the latter has been set to the half of the lowest reference values to model a more effective viral infection. We also consider $R_{v}=R_{u}$ in the initial conditions to model a wide initial infection. The blue dots show the single results greater than zero. The white dots show the median, which is zero in both cases. (c)–(d) Infection and tumour control probabilities for the two agent-based models (undirected movement in solid blue and pressure-driven movement in solid red) and their corresponding continuous counterparts (same colours, with dotted lines).
Continuum models cannot describe stochastic extinctions, but still provide meaningful information. In particular, the minimal cell number of the solution of Eq. (U3) is lower than the analogous quantity for Eq. (P3), in line with the greater effectiveness of the therapy in the individual counterpart of Eq. (U3). On the other hand, the continuum models do not predict the higher growth in the pressure-driven model: deterministic dynamics become less relevant in the regrowth, as the initial state for this last phase is significantly affected by the result of the previous stochastic events. Continuum models also provide a good estimate of the tumour control probability (TCP), as explained in Ref. [4]; nevertheless, we observed that the continuous TCP in both cases underestimates the corresponding quantity for the agent-based model. The same approach allows to correctly estimate the infection control probability (ICP).
@@
Our results show that, at least in principle, the virotherapy may cause a complete tumour remission if the viral infection is extremely effective; it remains an open question how to translate this in a clinical settings. We remark that for such low cell densities the immune clearance of the surviving cells could facilitate tumour extinction; on the other hand, an early immune response may hinder the infection and lead to an overall decrease of the therapy effectiveness, as discussed in Ref. [47].
@@ 4 Conclusions
In this work, we have investigated the behaviour of several models of oncolytic virotherapy when viral dynamics is explicitly formulated in the equations. A comparison with known and new treatment scenarios for systems that do not involve such mathematical modelling has highlighted some valuable ideas and also improved our knowledge of the role of motility of single and collective tumour cell(s) when a virus is present.
For explicit models, the rules governing the spatial movement of viruses appear to affect the infection more significantly than the rules governing tumour cell movement. Nonetheless, cell movement does still retain its influence on tumour invasion, especially when movement is constrained: only pressure-driven equations describe a slower invasion for low tumour cell numbers. This is a relevant aspect in the overall success or failure of the therapy and bears clinical consequences that are important when analysing clinical results. Overall, undirected movement (with or without viral diffusion) and pressure-driven movement without viral diffusion exhibit similar behaviour in regards to the spread of the infection throughout a tumour. The main difference among all models is the speed of the infections and, when this is crucial for applications, our result suggest that the viral population should be openly included in the formulation. An additional difference that has therapeutic value is the appearance of oscillations between viral and tumour populations, which is not feature shared by all models. For instance, a pressure-driven tumour not explicitly infected by a viral population (or with a low viral diffusion coefficient) exhibits a very different behaviour than a model with unhindered tumour diffusion. Notably, the presence of a viral population can also lead to persistent oscillations, which, as we have seen, can significantly affect the overall outcome in some cases. This could be very valuable when considering, for example, the effect of viral vectors on outcomes for some chronic blood tumours, which have been known for many years to exhibit semi-regular oscillatory regimes [66, 14, 57].
@@
In conclusion, it is important to remark that, for pressure-driven models, cells’ displacement alone cannot spread viral infection and central infections remain localised. As previously observed, even wide infections may be stopped by stochastic events [46], but the situation significantly changes when viral particles are allowed to freely move around the tumour and cell-to-cell contact becomes less relevant for the infection. In that case, our theoretical and computational results have explained that this infection is qualitatively similar to the case of undirected cell movement. We can conclude that pressure-driven movement allows us to describe both the situations of constrained and unconstrained cellular and viral movements, while such constraints cannot be modelled with undirected movement. The increased realism of pressure-driven movement, though, comes at the cost of a lower degree of theoretical understanding of the models.
@@
Appendix A Speed of the travelling waves (following Ref. [4])
@@ A.1 Uninfected cells
$\partial_{\hat{t}}C=\partial_{\hat{x}}(C\partial_{\hat{x}}C)+(1-C)C$
Results analogous to the previous case still hold, but this time the minimum speed is $c^{*}\coloneqq 1/\sqrt{2}$ [3, 48]. Its dimensional value is $\sqrt{Dp/2}$. It is straight forward to check that we recover the same speed value by multiplying the diffusion coefficient by a factor of $8$: this is indeed the procedure that we follow whenever we compare models with different rules of movement, so that we avoid any bias due to the different propagation speed.

### v2

@@ 3.4 Significant treatment success: emergence of oscillations
As we anticipated, it is well known that solutions of Eq. (U3) may exhibit persistent oscillations for some parameter values; this is in line with the fact that the non-spatial model of Eq. (2.6) presents a Hopf bifurcation. Fig. 6a shows a codimension one bifurcation plot in $\alpha$, whereas Fig. 6b-c shows the time series for three different values of $\alpha$. We remark that a similar behaviour is also observed, for example, as $\beta$} increases or as $q_{v}$ decreases. The one-dimensional setting is not ideal for studying oscillations, as local tumour extinctions due to stochasticity are very common: indeed, the typical behaviour observed in such parameter ranges is analogous to the one described below in relation to Fig. 10. We therefore concentrate our attention to two spatial dimensions: the higher number of cells involved allow us to describe oscillations, as well as stochastic-driven partial and complete eradication.
Figure 6: (a) One parameter bifurcation plot in $\alpha$ of Eq. (2.2), with other parameters as in Table 1 and $q_{v}=0.167\,$h-1. The green dots show the maximum and minimum values of $u$ for the stable limit cycle oscillations. The solid lines show the equilibrium value of $u$; the line is red if the equilibrium is stable and black if it is unstable. Hopf bifurcations are denoted by HB. Observe that for low values of $\alpha$, the infection-free equilibrium close to carrying capacity is stable. (b-c) Numerical simulation of Eq. (2.5) with the parameters as in Table 1 and different values of the viral burst size $\alpha$. The solid blue lines represent uninfected cells, the solid red lines infected cells and the dotted purple lines represent viral density multiplied by $q_{v}/(\alpha q)$; the two latter quantities superimpose almost exactly. The oscillations become wider as $\alpha$ increases, in accordance with the bifurcation diagram of panel (a).
@@
Fig. 7, along with the video accompanying it (see electronic supplementary material S7), shows an overall excellent quantitative agreement between numerical solutions of the system of PDEs (U3) and the average over five numerical simulations of the corresponding agent-based model in two spatial dimensions. After the initial infection reaches the tumour boundary, the number of cancer cells decreases significantly; as a consequence, stochasticity becomes relevant and, in some regions, cells go extinct. Fig. 7a shows the inhomogeneous pattern of uninfected cells caused by this. Similar patterns would also arise in the two-dimensional non-radial continuum model under stochastic perturbations. As time elapses, the spatial inhomogeneities quickly disappear, and both models show similar persistent oscillations in the tumour centre around the equilibrium values (coherently with what the bifurcation diagram shows). A slight delay in the oscillations is the main difference between the two modelling approaches: this is easy to see from the total number of cells, depicted in Fig. 9a. The invasion speed of the tumour in the surrounding area is still $2\sqrt{Dp}$ and the uninfected cell density at the front is approximately given by Eq. (3.2), despite some fluctuations. It is interesting to note that highly dynamic patterns may emerge when viral injections at multiple sites are considered [baabdulla23]: this can be easily checked through VisualPDE [walker23], using the simulation available at the link https://visualpde.com/sim/?mini=raZybpXG.
Fig. 9a shows the comparison with the solution of Eq. (U2) to remark once again that oscillations cannot be obtained in the two-equations model even for high infection rates. Furthermore, Fig. 9b shows that oscillations are almost non-existent for lower viral diffusion, resembling the behaviour of Eq. (U2); this confirms once again the excellent quality of the quasi-steady assumption as spatial viral dynamics become negligible and the reactions involving the virus increase their rate.
@@
Figure 9: Comparison between the sum of tumour cells in different scenarios. The parameters employed are the ones given in Table 1, with $\alpha=3500\;$viruses/cells and $q_{v}=1.67\times 10^{-1}\;$h-1; $D_{v}$ assumes different values in panels (b) and (c). Solid lines represent the average of five agent-based simulations in the case of undirected cell movement (yellow and red) and pressure-driven cell movement (purple and green). All the other lines represent the sum of cells obtained from continuous models. The dashed blue lines refer to the solution of Eq. (U2); they are in a different colour than the other PDE solutions, to highlight the fact that it is related to a two-equations model, which does not exhibit persistent oscillations. The black lines refer to the solutions of either Eq. (U3) or Eq. (P3): in panels (b) and (c) we show for simplicity only the case of lower viral diffusion, as the reference cases are already shown in panel (a) and their qualitative behaviour is similar to the discrete models.
@@
Fig. 8a, along with the video accompanying it (see electronic supplementary material S8), shows that, as in the previous case, the number of cancer cells decreases significantly after the initial infection reaches the tumour boundary in both models and, in some regions, cells go extinct, originating inhomogeneous patterns. However, the subsequent regrowth does not allow a fast spatial spread due to the slow movement that cells experience when the pressure is low. Fig. 8b indeed shows that uninfected cells take a long time in the agent-based model to move from the hollow ring into the central region; on the other hand, in the case of the PDE, the repopulation is much faster: there is never a complete extinction in the centre and the increase of cells’ number in some areas is caused by regrowth rather than movement. This difference is particularly evident from the second peak in Fig. 9a. The difference between the two models tends to disappear as time goes on, but Fig. 8c still shows some dissimilarities.
Both models have an invasion speed much lower than $\sqrt{Dp/2}$: indeed, the nonlinear diffusion is affected by the scale of the cell density, as explained in Appendix A. In our case, the oscillatory behaviour of the system does not allow to easily identify a unique, invariant speed. Fig. 8 shows that we can obtain a good approximation by setting $\bar{u}=1800\;$cells/mm2, which corresponds to an average of the peak over time. The change in the speed in presence of effective infections is a significant difference from the case of undirected cell movement and, as a result, the overall cell numbers involved are lower in the second case (see Fig. 9a). It is also interesting to remark that inhomogeneous infection patterns may result in an anisotropic spread of the tumour in the surrounding environment: we refer the interested reader to the VisualPDE simulation available at the link https://visualpde.com/sim/?mini=e0LGK0st.
@@
The oscillations analysed so far are able to drive the tumour at low densities, in which stochastic fluctuations may cause the death of all the agents in some regions; nevertheless, oscillations are not wide enough to lead the whole tumour close to extinction. It is important to see if it is possible to increase the effectiveness of the therapy and have the possibility of a full tumour remission. We therefore decrease the value of $q_{v}$ to half of the reference value, so that the oscillations can reach lower cell densities. We also consider an initial wide viral infection (i.e., $R_{v}=R_{u}$) to study the effect of a higher initial viral load. Fig. 10a-b summarise the cell number of one hundred simulations for both cell movement at time t=1500t=1500\;h. In both cases, the wide majority of the simulations shows the extinction of the tumour; the case of undirected cell movement appears more successful, with only eight simulations with a surviving tumour (in contrast to nineteen in the other case) and an overall lower cell density at the final time. It is convenient to track the distribution over time of the extinctions of the agent-based model through the tumour control probability, defined as
[…]
Fig. 10d shows the eradication of the infection in all the simulations performed. This explains the binary outcome that we observe: either all cancer cells become infected and eventually die out during the first oscillation, or a few cells avoid the infection and give rise to a recrudescence of the cancer.
@@
Figure 10: (a)-(b) Violin plots of the final cell number at time t=1500t=1500\;h obtained from one hundred simulations of the agent-based models with the parameters given in Table 1, $\alpha=3500\;$viruses/cells and $q_{v}=8.33\times 10^{-2}\;$h-1; the latter has been set to the half of the lowest reference values to model a more effective viral infection. We also consider $R_{v}=R_{u}$ in the initial conditions to model a wide initial infection. The blue dots show the single results greater than zero. The white dots show the median, which is zero in both cases. (c)–(d) Tumour and infection control probabilities for the two agent-based models from Eqs. (3.3) and Eq. (3.4) (undirected movement in solid blue and pressure-driven movement in solid red) and their corresponding continuous counterparts via the Poissonian assumption from Eqs. (3.5) and Eq. (3.6) (same colours, with dotted lines).
Continuum models cannot describe stochastic extinctions, but still provide meaningful information. In particular, the minimal cell number of the solution of Eq. (U3) is lower than the analogous quantity for Eq. (P3), in line with the greater effectiveness of the therapy in the individual counterpart of Eq. (U3). A more quantitative comparison can be made through the Poissonian tumour control probability
$\text{TCP}(t)\coloneqq\exp\Bigl[-2\pi\int_{0}^{R}u(t,r)\,\mathrm{d}r\Bigr]$ (3.5)
which is based on the assumption that the number of surviving uninfected cells follows a Poisson distribution. In Ref. [gong13], the Poissonian TCP defined as above was compared with more complex expressions and it was found to perform equally well in the case of radiotherapy. Fig. 10c shows an excellent quantitative agreement at early times between the expressions of Eqs. (3.3) and (3.5). The agreement is clearly lost as the continuum model shows the recrudescence of the tumour. It is also interesting to observe that the maximum value of the Poissonian TCP is slightly lower than the asymptotic value of the discrete TCP in both cases. A possible explanation for this discrepancy is that localised cell clusters are more effectively infected than wider, spatially homogeneous populations with a lower density; in the agent-based models, stochastic fluctuations during the infection naturally generate such small clusters, whereas the continuum model represents only averages and therefore cannot reproduce this effect. Nevertheless, the Poissonian TCP still provides a meaningful insight on the underlying stochastic model.
We can follow the same approach to define the Poissonian infection control probability
$\text{ICP}(t)\coloneqq\exp\Bigl[-2\pi\int_{0}^{R}i(t,r)\,\mathrm{d}r\Bigr]$ (3.6)
Fig. 10d shows an excellent quantitative agreement between the expressions of Eqs. (3.4) and (3.6), meaning that the discrepancies mentioned above do not affect the dynamics of infected cells.
@@
Our results show that, at least in principle, the virotherapy may cause a complete tumour remission if the viral infection is extremely effective; this is interesting, but it remains an open question how to translate this in a clinical settings. We remark that, for such low cell densities, the immune clearance of the surviving cells could facilitate tumour extinction; on the other hand, an early immune response may hinder the infection and lead to an overall decrease of the therapy effectiveness, as discussed in Ref. [morselli24].
@@ 4 Conclusions
Undirected cell movement Pressure-driven cell movement
Quasi-steady virus Continuum model: Eq. (U2) Continuum model: Eq. (P2)
Comprehensive theoretical understanding Partial theoretical understanding
Constant tumour invasion speed Slower tumour invasion at low cell densities
Central infections infect all the tumour Central infections cannot spread
No significant stochastic effects Significant stochastic effects
No oscillatory behaviour No oscillatory behaviour
Explicit virus Continuum model: Eq. (U3) Continuum model: Eq. (P3)
Comprehensive theoretical understanding Partial theoretical understanding
Constant tumour invasion speed Slower tumour invasion at low cell densities
Central infections infect all the tumour (with a speed that may be different than in case of Eq. (U2)) Central infections only spread with high $D_{v}$ (with a speed that may be different than in case of Eq. (U2))
Significant stochastic effects in presence of oscillations that lead to partial or complete extinction Significant stochastic effects for low $D_{v}$ and in presence of oscillations that lead to partial or complete extinction
Table 2: Comparison of tumour invasion models with implicit vs explicit virus formulation and different cell movement assumptions.
In this work, we have investigated the behaviour of several models of oncolytic virotherapy when viral dynamics is explicitly formulated in the equations. A comparison with known and new treatment scenarios for systems that do not involve such mathematical modelling has highlighted some valuable ideas and also improved our knowledge of the role of motility of single and collective tumour cell(s) when a virus is present. The results are summarised in Table 2.
For explicit models, the rules governing the spatial movement of viruses appear to affect the infection more significantly than the rules governing tumour cell movement. Nonetheless, cell movement does still retain its influence on tumour invasion, especially when movement is constrained: only pressure-driven equations describe a slower invasion for low tumour cell numbers. This is a relevant aspect in the overall success or failure of the therapy and bears clinical consequences that are important when analysing patient’s results. Overall, undirected movement (with or without viral diffusion) and pressure-driven movement with viral diffusion exhibit similar behaviour in regards to the spread of the infection throughout a tumour. The main difference among all models is the speed of the infections and, when this is crucial for applications, our result suggest that the viral population should be openly included in the formulation. An additional difference that has therapeutic value is the appearance of oscillations between viral and tumour populations, which is not a feature shared by all models. For instance, a pressure-driven tumour not explicitly infected by a viral population (or with a low viral diffusion coefficient) exhibits a very different behaviour than a model with unhindered tumour diffusion. Notably, the presence of a viral population can also lead to persistent oscillations, which, as we have seen, can significantly affect the overall outcome in some cases. This could be very valuable when considering, for example, the effect of viral vectors on outcomes for some chronic blood tumours, which have been known for many years to exhibit semi-regular oscillatory regimes [CHIKKAPPA1980Blood, RODRIGUEZ1976AMJ, Vodopick1972NEJM].
@@
In conclusion, it is important to remark that, for pressure-driven models, cells’ displacement alone cannot spread viral infection and central infections remain localised. As previously observed, even wide infections may be stopped by stochastic events [morselli23], but the situation significantly changes when viral particles are allowed to freely move around the tumour and cell-to-cell contact becomes less relevant for the infection. In that case, our theoretical and computational results have explained that this infection is qualitatively similar to the case of undirected cell movement. We can conclude that pressure-driven movement allows us to describe both the situations of constrained and unconstrained cellular and viral movements, while such constraints cannot be modelled with undirected movement. The increased realism of pressure-driven movement, though, comes at the cost of a lower degree of theoretical understanding of the models.
@@
Appendix A Speed of the travelling waves
@@ A.1 Uninfected cells
$\partial_{\hat{t}}C=\partial_{\hat{x}}(C\partial_{\hat{x}}C)+(1-C)C$ (A.2)
Results analogous to the previous case still hold, but this time the minimum speed is $c^{*}\coloneqq 1/\sqrt{2}$ [aronson80, newman80]. Its dimensional value is $\sqrt{Dp/2}$. It is straightforward to check that we recover the same speed value as in Eq. (A.1) by multiplying the diffusion coefficient by a factor of $8$: this is indeed the procedure that we follow whenever we compare models with different rules of movement, so that we avoid any bias due to the different propagation speed.